Need some help understanding oxidation values

I am needing to better understand the oxidation numbers assigned to compounds. On one hand, the oxidation number of a neutral compound is 0, such as H2O. But, then I have another rule in my notes that states "In a binary compound, the oxidation number is equal to the charge of the element with the greatest electronegativity", an example being PCl3, which has an oxidation number of -1. So how do I make sense of this? It seems conflicting to me, why isn't H2O's oxidation value -2? Why isn't PCl3's 0? Can someone please break it down for me and explain how each applies. Thank you
Answer #1

First of all the oxidation number is defined for elements in compound.for example

In H2O , we calculate oxidation number of H and O ,not H2.

As you saying zero for some compound.

It is that the algebric sum of oxidation number of all atoms in a neutral compound is zero. For example in H2O

(Oxidation number of H) + (oxidation number of O) = 0

2×(+1) + 1×(-2) = 0 (as there are two H each has oxidation number +1 and one O has oxidation number -2)

And within compound higher electronegative element have negative charge.

In PCl3 also ,sum of oxidation number of all atoms in compound is zero.

1×(+3) + 3×(-1) = 0

But for charged compound,the sum of oxidation number equal to charge on compound, for example

SO​​​​​42-

Oxidation number of O + oxidation number of S = -2.

In both example as you mentioned overall neutral compound has sum of oxidation number is 0